# Yang-Mills extension of the Loop Quantum Gravity-corrected Maxwell equations

### v1

@@ III The Noether Procedure as our path towards the non-Abelian scenario
$\begin{split}\vec{j}_{h}=-f_{hij}\{A_{\gamma}[(\vec{E}_{i}\cdot\phi_{j})-(\vec{B}_{i}\times\vec{A}_{j})]-\theta_{3}l_{P}^{2}[(\vec{B}_{i}\times\\ \nabla^{2}\vec{A}_{j})-(\nabla^{2}\vec{B}_{i}\times\vec{A}_{j})-(\nabla^{2}\vec{E}_{i}\cdot\phi_{j})-(\vec{E}_{i}\cdot\nabla^{2}\phi_{j})]\\ -\theta_{8}l_{P}[2(\nabla^{2}\vec{A}_{i}\times\vec{A}_{j})-(\vec{E}_{i}\times\vec{\dot{A}}_{j})-(\nabla\times\vec{E}_{i})\phi_{j}]\\ -4\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(B^{2}\cdot\vec{B}_{i}\times\vec{A}_{j})\},\end{split}$ (18)
@@
$\begin{split}\vec{E}_{h}=A_{\gamma}[-\nabla\cdot\phi_{h}-\vec{\dot{A}}_{h}+gf_{hij}\phi_{i}\vec{A}_{j}]-\frac{3g}{2}\theta_{3}l_{P}^{2}f_{hij}[\vec{A}_{i}\\ (\nabla^{2}\phi_{j})+(\nabla^{2}\vec{A}_{i})\cdot\phi_{j}]+\frac{g}{2}\theta_{8}l_{P}f_{hij}(\vec{A}_{i}\times\vec{\dot{A}}_{j}),\end{split}$ (25)
@@
$\begin{split}\vec{B}_{h}=A_{\gamma}[(\nabla\times\vec{A}_{h})+\frac{g}{2}f_{hij}(\vec{A}_{i}\times\vec{A}_{j})]-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot\\ [\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-[\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})]]+\frac{g}{2}\theta_{8}l_{P}f_{hij}\cdot\\ [(\vec{\dot{A}}_{i})\cdot\phi_{j}]-6g\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}.\end{split}$ (26)
@@
$\frac{\partial H}{\partial\vec{E}_{h}}=-\frac{\partial\vec{A}_{h}}{\partial t},$ (27)
@@
$\frac{\partial H}{\partial\vec{A}_{h}}=\frac{\partial\vec{E}_{h}}{\partial t}.$ (28)
@@
$\begin{split}A_{\gamma}[\nabla\cdot\vec{E}_{h}+gf_{hij}\vec{E}_{i}\cdot\vec{A}_{j}]+\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}[2(\nabla^{2}\vec{E}_{i})\cdot\vec{A}_{j}+\vec{E}_{i}\cdot(\nabla^{2}\vec{A}_{j})+\nabla^{2}(\vec{A}_{i}\cdot\vec{E}_{j})-\\ 2\nabla[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]]+\frac{g}{2}\theta_{8}l_{P}f_{hij}\vec{A}_{i}\cdot(\nabla\times\vec{E}_{j})=0,\end{split}$ (29)
$\begin{split}A_{\gamma}[\nabla\times\vec{B}_{h}-gf_{hij}\phi_{i}(\vec{\dot{A}}_{j})-gf_{hij}\phi_{i}\vec{E}_{j}+gf_{hij}(\vec{B}_{i}\times\vec{A}_{j})]+\theta_{3}l_{P}^{2}\{2\nabla^{2}(\nabla\times\vec{B}_{h})-gf_{hij}[\\ [(\nabla^{2}\vec{E}_{i})\phi_{j}+\frac{1}{2}\nabla^{2}(\phi_{i}\vec{E}_{j})+[(\nabla^{2}\vec{B}_{i})\times\vec{A}_{j}]-\frac{1}{2}\phi_{i}\nabla^{2}(\vec{\dot{A}}_{j})+\frac{1}{2}\vec{E}_{i}(\nabla^{2}\phi_{j})-\frac{1}{2}(\vec{\dot{A}}_{i})\cdot(\nabla^{2}\phi_{j})-\\ \frac{1}{2}\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})-\frac{1}{2}[\vec{B}_{i}\times(\nabla^{2}\vec{A}_{j})]-\nabla^{2}[\phi_{i}(\vec{\dot{A}}_{j})]]\}-\theta_{8}l_{P}\{2\nabla^{2}\vec{B}_{h}-\frac{g}{2}f_{hij}[\partial_{t}(\vec{A}_{i}\times\vec{A}_{j})-\\ [(\vec{\dot{E}}_{i})\times\vec{A}_{j}]-2\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})+(\nabla\times\vec{E}_{i})\phi_{j}]\}+4\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times(B^{2}\vec{B}_{h})=\frac{\partial\vec{E}_{h}}{\partial t},\end{split}$ (30)
$\begin{split}A_{\gamma}[(\nabla\cdot\vec{B}_{h})+gf_{hij}\vec{A}_{i}\cdot\vec{B}_{j}]-g\theta_{3}l_{P}^{2}f_{hij}[\nabla^{2}(\vec{A}_{i}\cdot\vec{B}_{j})-\vec{A}_{i}\cdot(\nabla^{2}\vec{B}_{j})]+\frac{g}{2}\theta_{8}l_{P}f_{hij}\nabla\cdot[(\vec{\dot{A}}_{i})\cdot\phi_{j}]\\ -12g\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}f_{hij}\vec{A}_{i}\cdot(\vec{B}_{j}\cdot B^{2})=0,\end{split}$ (31)
$\begin{split}A_{\gamma}[(\nabla\times\vec{E}_{h})-gf_{hij}\phi_{i}\vec{B}_{j}+gf_{hij}(\vec{A}_{i}\times\vec{E}_{j})]+\theta_{3}l_{P}^{2}\nabla^{2}(\nabla\times\vec{E}_{h})-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}[\vec{B}_{i}\cdot(\nabla^{2}\phi_{j})\\ -(\nabla^{2}\vec{E}_{i}\times\vec{A}_{j})]-\theta_{8}l_{P}[\nabla^{2}\vec{E}_{h}+gf_{hij}\nabla\cdot(\vec{E}_{i}\cdot\vec{A}_{j})-\frac{g}{2}f_{hij}\nabla\times(\vec{A}_{i}\times(\vec{\dot{A}}_{j}))]=-\frac{\partial\vec{B}_{h}}{\partial t}.\end{split}$ (32)
@@ V Acknowledgements
G.L.L.W.L. acknowledges financial support of the Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/202.437/2024. We are also grateful to P. A. Lima Mourão for valuable discussions on the LQG-corrected Maxwell equations in presence of a background magnetic field.

### v2

@@ III The Noether Procedure as our path towards the non-Abelian scenario
$\begin{split}\vec{j}_{h}=-f_{hij}\{A_{\gamma}[(\vec{E}_{i}\cdot\phi_{j})-(\vec{B}_{i}\times\vec{A}_{j})]-\theta_{3}l_{P}^{2}[(\vec{B}_{i}\times\\ \nabla^{2}\vec{A}_{j})-(\nabla^{2}\vec{B}_{i}\times\vec{A}_{j})-(\nabla^{2}\vec{E}_{i}\cdot\phi_{j})-(\vec{E}_{i}\cdot\nabla^{2}\phi_{j})]\\ -\theta_{8}l_{P}[2(\nabla^{2}\vec{A}_{i}\times\vec{A}_{j})-(\vec{E}_{i}\times\vec{\dot{A}}_{j})-(\nabla\times\vec{E}_{i})\phi_{j}]\\ +4\theta_{4}\mathcal{L}^{2}l_{P}^{2}\left(\frac{\mathcal{L}}{l_{p}}\right)^{2\Upsilon}(B^{2}\cdot\vec{B}_{i}\times\vec{A}_{j})\},\end{split}$ (18)
@@
$\begin{split}\vec{E}_{h}=A_{\gamma}[-\nabla\cdot\phi_{h}-\vec{\dot{A}}_{h}+gf_{hij}\phi_{i}\vec{A}_{j}]-\frac{3g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot\\ [\vec{A}_{i}(\nabla^{2}\phi_{j})+(\nabla^{2}\vec{A}_{i})\cdot\phi_{j}]+\theta_{3}l_{P}^{2}f_{hij}[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]\\ +\frac{g}{2}\theta_{8}l_{P}f_{hij}(\vec{A}_{i}\times\vec{\dot{A}}_{j}),\end{split}$ (25)
@@
$\begin{split}\vec{B}_{h}=A_{\gamma}[(\nabla\times\vec{A}_{h})+\frac{g}{2}f_{hij}(\vec{A}_{i}\times\vec{A}_{j})]+2l_{P}^{2}\theta_{3}\nabla^{2}\vec{B}_{h}\\ -\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\cdot[\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-[\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})]]+\\ \frac{g}{2}\theta_{8}l_{P}f_{hij}[(\vec{\dot{A}}_{i})\cdot\phi_{j}]+2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}[2(\vec{B}^{2}\vec{B}_{h})\\ +3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}].\end{split}$ (26)
@@
$\frac{\partial H_{LQG}^{2}}{\partial\vec{E}_{h}}=-\frac{\partial\vec{A}_{h}}{\partial t},$ (27)
@@
$\frac{\partial H_{LQG}^{2}}{\partial\vec{A}_{h}}+\frac{\partial H_{LQG}^{2}}{\partial(\nabla\times\vec{A}_{h})}=\frac{\partial\vec{E}_{h}}{\partial t}.$ (28)
@@
$2A_{\gamma}[\nabla\cdot\vec{E}_{h}+gf_{hij}\vec{E}_{i}\cdot\vec{A}_{j}]+g\theta_{3}l_{P}^{2}f_{hij}[2(\nabla^{2}\vec{E}_{i})\cdot\vec{A}_{j}+\vec{E}_{i}\cdot(\nabla^{2}\vec{A}_{j})+\nabla^{2}(\vec{A}_{i}\cdot\vec{E}_{j})-$
 $\;2\nabla[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]]+g\theta_{8}l_{P}f_{hij}[\nabla\cdot(\vec{A}_{i}\times\vec{\dot{A}}_{j})+\vec{A}_{i}\cdot(\nabla\times\vec{E}_{j})]=0,$ (29)
$A_{\gamma}[\nabla\times\vec{B}_{h}-gf_{hij}\phi_{i}(\vec{\dot{A}}_{j})-gf_{hij}\phi_{i}\vec{E}_{j}+gf_{hij}(\vec{B}_{i}\times\vec{A}_{j})]+\theta_{3}l_{P}^{2}\{2\nabla^{2}(\nabla\times\vec{B}_{h})-gf_{hij}[(\nabla^{2}\vec{E}_{i})\phi_{j}+$
$\;\frac{1}{2}\nabla^{2}(\phi_{i}\vec{E}_{j})+[(\nabla^{2}\vec{B}_{i})\times\vec{A}_{j}]-\frac{1}{2}\phi_{i}\nabla^{2}(\vec{\dot{A}}_{j})+\frac{1}{2}\vec{E}_{i}(\nabla^{2}\phi_{j})-\frac{1}{2}(\vec{\dot{A}}_{i})\cdot(\nabla^{2}\phi_{j})-\frac{1}{2}\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})-$
$\;\frac{1}{2}[\vec{B}_{i}\times(\nabla^{2}\vec{A}_{j})]-\nabla^{2}[\phi_{i}(\vec{\dot{A}}_{j})]]\}-\theta_{8}l_{P}\{2\nabla^{2}\vec{B}_{h}-\frac{g}{2}f_{hij}[\partial_{t}(\vec{A}_{i}\times\vec{A}_{j})-[(\vec{\dot{E}}_{i})\times\vec{A}_{j}]-2\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-$
$\;\nabla[\nabla\cdot(\vec{A}_{i}\times\vec{A}_{j})]+(\nabla\times\vec{E}_{i})\phi_{j}]\}+2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times\left[2(\vec{B}^{2}\vec{B}_{h})+3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}\right]=\frac{\partial\vec{E}_{h}}{\partial t},$ (30)
$A_{\gamma}[(\nabla\cdot\vec{B}_{h})+gf_{hij}\vec{A}_{i}\cdot\vec{B}_{j}]-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\nabla\cdot[(\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})-3\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})]+\frac{g}{2}\theta_{8}l_{P}f_{hij}\nabla\cdot[(\vec{\dot{A}}_{i})\cdot\phi_{j}]+$
 $\;9g\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}f_{hij}\nabla\cdot[(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}]=0,$ (31)
$A_{\gamma}[(\nabla\times\vec{E}_{h})-gf_{hij}\phi_{i}\vec{B}_{j}+gf_{hij}(\vec{A}_{i}\times\vec{E}_{j})]+\theta_{3}l_{P}^{2}\nabla^{2}(\nabla\times\vec{E}_{h})-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}[\vec{B}_{i}\cdot(\nabla^{2}\phi_{j})-$
$\;(\nabla^{2}\vec{E}_{i}\times\vec{A}_{j})]-\theta_{8}l_{P}[\nabla^{2}\vec{E}_{h}+gf_{hij}\nabla\cdot(\vec{E}_{i}\cdot\vec{A}_{j})-\frac{g}{2}f_{hij}\nabla\times(\vec{A}_{i}\times(\vec{\dot{A}}_{j}))]=-\frac{\partial\vec{B}_{h}}{\partial t}.$ (32)
@@ V Acknowledgements
G.L.L.W.L. acknowledges financial support of the
Fundação Carlos Chagas Filho de Amparo à Pesquisa
do Estado do Rio de Janeiro (FAPERJ), Grant No. E-
26/202.437/2024. We are also grateful to P. A. Lima
Mourão for valuable discussions on the LQG-corrected
Maxwell equations in presence of a background magnetic field.